O 6.95 Question 26 Calculate the pH of a buffer solution made of 0.9 M NH3 and 0.3 M NH4Cl? Kb= 1.8x10^-5 (1 point ) * O 2.88 O 4.27 5.9 O 9.73 Question 27 Which of the following dissociates into ions the most in
Answer #1

[NH3]    = 0.9M

[NH4Cl]   = 0.3M

PKb   = -logKb

          = -log1.8*10^-5

           = 4.74

POH   = Pkb + log[NH4Cl]/[NH3]

           = 4.75 + log0.3/0.9

           = 4.75-0.477

          = 4.273

PH   = 14-POH

        = 14-4.273

         = 9.73

9.73 >>>>>answer